# Robust Max Statistics for High-Dimensional Inference

### v1

@@ 4.2 Functional data
Discussion of results. In Figures 3(a)-3(d), the rejection rate curves for the proposed method are labeled by PM, and the corresponding curves based on the intervals proposed in (Fan:2023, ) are labeled HL. Recall that $h=0$ under the null hypothesis $\mathsf{H}_{0}$, and so the value of a curve at $h=0$ represents the empirical level. It is clear that in all four panels of Figure 3, the empirical levels of both methods closely match the nominal level of $\alpha=5\%$, marked with a dashed horizontal line.
However, the methods differ markedly in terms of power, which is represented by the values of the curves at $h>0$. In the three settings based on the stock data of Apple, Nvidia, and JP Morgan, the power of PM increases steadily with $h$, whereas the power of HL stays relatively flat. In the setting based on Moderna stock data, the two methods are more competitive, but even here, the power of PM is still noticeably higher for most values of $h$. The power advanatage of PM is understandable in light of the characteristics of the simultaneous confidence intervals ^$\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\beta_{1},\dots,\beta_{p}$. (Recall that both methods reject the null hypothesis $\mathsf{H}_{0}:h=0$ whenever any of their associated intervals exclude 0.) As was observed in Section 4.1, the intervals produced by PM tend to be tighter than those produced by HL, and tighter intervals make it easier to exclude 0, resulting in higher power.
@@
4.3 Application to financial data
In this subsection, we study an application that uses historical log return curves in order to screen candidate pairs of stocks for a strategy known as pairs trading. In addition to its interest from the viewpoints of finance and functional data analysis, this application illustrates two effects that are related to our proposed methodology: heavy tails and variance decay.
Background on pairs trading. Pairs trading is a well-established strategy that takes advantage of highly correlated pairs of stocks when their relative performance deviates from typical behavior (pairs, ). When these situations occur, the strategy involves purchasing the stock that is undeperforming, and selling short the stock that is overperforming—with the idea being that both trades will become profitable when the relative performance reverts to its mean. One of the main reasons for interest in pairs trading is that even though the individual performances of the stocks may be hard to predict, the relative performance may still be rather stable, which enhances the reliability of the strategy.
Below, we apply the method proposed in Section 2 to study a basic question that is essential to any pairs trading strategy. Starting from a collection of candidate pairs of stocks, we seek to identify which pairs have relative performance that is unstable. In other words, we seek to identify which pairs should be discarded from further consideration, because they are not well suited to pairs trading. Hence, this problem can be viewed as part of a preliminary filtering process that is done before any trading takes place.
Problem formulation. For a given pair of stocks, we measure their relative performance in terms of the difference $D(t)$ of their log returns (ruppertstatistics, ), which is viewed as a random function of time $t\in[0,1]$. As a way of preparing an approximate sample of realizations of $D(t)$, we used historical stock data, as discussed later in this subsection. To formalize the question of whether the relative performance is stable or unstable, we consider the hypothesis testing problem
[…]
Thus, if $\mathsf{H}_{0}$ cannot be rejected, then the stocks underlying $D(t)$ are plausible candidates for pairs trading. It is also worth noting that a variety of other approaches have been proposed for selecting trading pairs (pairs, , Ch. 5-6), and the formulation above is only intended to serve as an illustration of our proposed methodology. More generally, the problem of testing whether or not a mean function is constant is of fundamental interest in other applications involving functional data.
Although the null hypothesis $\mathsf{H}_{0}$ might seem easy to reject at first sight, it should be emphasized that $\mathsf{H}_{0}$ is often difficult to reject for three reasons. First, the value of $\mathbf{E}(D(0))$ under $\mathsf{H}_{0}$ is unknown and unrestricted, which means that $\mathsf{H}_{0}$ allows for many possible values of $\mathbf{E}(D(0))$. Consequently, it is more difficult to reject $\mathsf{H}_{0}$ than a simpler condition such as ‘$\mathbf{E}(D(t))=c_{0}$ for all $t\in[0,1]$’, where $c_{0}$ is a particular value that is specified by the user. Second, as is commonly done in pairs trading, we used data for pairs of stocks involving similar businesses. Hence, this predisposes $\mathbf{E}(D(t))$ to be more stable than it would be for an arbitrarily chosen pair of stocks. Third, the log return of a generic stock typically fluctuates near 0, which reflects the inherent difficulty of making profitable trades. Likewise, the difference of log returns $D(t)$ also tends to fluctuate near 0, making it difficult to reject $\mathsf{H}_{0}$.
Data preparation. Price data for 10 pairs of stocks (listed in Table 2) were collected from the Alpha Vantage database (alphavantage2024, ) during every minute of trading between March 1, 2024 and March 22, 2024, including pre-market and after-hours trading. The data were then divided into 30-minute intervals (normalized to unit length), and each interval was divided into time points $t_{0},t_{1},\dots,t_{30}$ spaced one minute apart. Letting $P_{1}(t_{j})$ and $P_{2}(t_{j})$ denote the prices for a pair of stocks at $t_{j}$, the difference of log returns was computed as $D(t_{j})=\log(P_{1}(t_{j})/P_{1}(t_{j-1}))-\log(P_{2}(t_{j})/P_{2}(t_{j-1}))$ for each $j=1,\dots,30$. In this way, we obtained one discretely observed realization of the function $D(t)$ over each 30-minute interval. To promote independence among these functional observations, we only retained them from every other 30-minute interval, ensuring that they are separated by gaps of 30 minutes. We also excluded functional observations that were obtained from intervals with missing data. The resulting set of functional observations are denoted $D_{1}(t),\dots,D_{n+m_{n}}$, and the values of $n$ and $m_{n}$ are shown in Table 2.
Testing procedure. To test whether or not the function $\mathbf{E}(D(t))$ is constant with respect to $t\in[0,1]$, we consider an expansion of the form $\mathbf{E}(D(t))=\sum_{j=1}^{\infty}\gamma_{j}\phi_{j}(t)$, where $\{\phi_{j}(t)\}_{j\geq 1}$ is again the Fourier cosine basis for $L^{2}[0,1]$. A convenient property of this basis is that $\mathsf{H}_{0}$ is equivalent to the condition that $\gamma_{j}=0$ for $j\geq 2$. (Note that $\phi_{1}(t)$ is a constant function, and so the first Fourier coefficient $\gamma_{1}$ may be still be non-zero under $\mathsf{H}_{0}$.) Similarly to our approach in Section 4.2, this leads us to consider the simultaneous hypothesis
$\small\mathsf{H}_{0,j}:\gamma_{j}=0\text{ \ \ \ for \ \ \ }j=2,\dots,p+1,$ (11)
where $p$ is chosen large enough such that $\gamma_{p+2},\gamma_{p+3},\dots$ are effectively negligible. As in Section 4.2, a procedure that controls the simultaneous type I error rate for $\mathsf{H}_{0,2},\dots,\mathsf{H}_{0,p+1}$ gives rise to one that controls the ordinary type I error rate for $\mathsf{H}_{0}$. Despite this similarity with Section 4.2, there is still a notable difference, since the first Fourier coefficient is irrelevant in the current context. This is because Section 4.2 dealt with testing the null that a function is identically zero, whereas here we are testing the null that a function is constant.
We applied the methodology from Section 2 to the random vectors $X_{1},\dots,X_{n+m_{n}}\in\mathbb{R}^{p}$, with $p=50$, where $X_{i}$ is defined to contain the coefficients of $D_{i}(t)$ with respect to $\phi_{2}(t),\dots,\phi_{p+1}(t)$, so that $\mathbf{E}(X_{i})=(\gamma_{2},\dots,\gamma_{p+1})$. Letting ^$\hat{\mathcal{I}}_{2},\dots,\hat{\mathcal{I}}_{p+1}$ denote the resulting confidence intervals for $\gamma_{2},\dots,\gamma_{p+1}$ with nominal simultaneous coverage probability of $1-\alpha$, we reject $\mathsf{H}_{0}$ with a nominal level of $\alpha$ whenever any of these intervals exclude 0.
Discussion of results. After testing 10 instances of $\mathsf{H}_{0}$ based on 10 pairs of stocks, we recorded the resulting p-values in the column of Table 2 labeled PM. For comparison, the column labeled HL includes p-values that result from replacing the simultaneous confidence intervals ^$\hat{\mathcal{I}}_{2},\dots,\hat{\mathcal{I}}_{p+1}$ in our testing procedure with the intervals proposed in (Fan:2023, ). For 8 of the 10 pairs of stocks, the table shows that PM produced p-values that are smaller than those for HL, often by a factor of 4 or more. Another way to compare the methods is to consider their ability to reject $\mathsf{H}_{0}$ at a 10% significance level. In this regard, PM is able to reject in 5 of the 10 cases, whereas HL is able to reject in none.
Table 2: Comparison of $p$-values produced by PM and HL
Company 1 Company 2 PM HL $n$ $m_{n}$
Apple Microsoft 0.118 0.318 230 20
Google Meta 0.372 0.290 210 20
Tesla General Motors 0.078 0.682 150 10
Amazon Walmart 0.046 0.296 180 10
Pfizer Moderna 0.188 0.886 150 10
Netflix Disney 0.096 0.384 180 20
Nvidia AMD 0.236 0.622 230 20
JPMorgan Chase Bank of America 0.062 0.282 140 10
Shopify eBay 0.458 0.270 100 10
UPS FedEx 0.060 0.704 100 10
To illustrate the difficulty of rejecting $\mathsf{H}_{0}$, Figures 4(a)-4(b) display representative log return curves over 100 minutes periods for the individual stocks in the (Apple, Microsoft) and (Google, Meta) pairs. Recalling that $D(t)$ is difference of the expected log return curves for each pair, these figures show that $\mathsf{H}_{0}$ is challenging to reject because the individual log return curves have few sustained movements away from 0.
To give additional context to these results, it is worth commenting on some of the empirical characteristics of the random vectors $X_{1},\dots,X_{n+m_{n}}$ of Fourier coefficients. As was discussed in Section 4.1, the HL method is appropriate for inference on the mean parameters $(\gamma_{2},\dots,\gamma_{p+1})=\mathbf{E}(X_{1})$ when the coordinates of $X_{1}$ have symmetric distributions around 0. In Figures 5(a)-5(d), we plot some representative histograms of these coordinate distributions for the pairs (Apple, Microsoft) and (Google, Meta), confirming that they are nearly symmetric. Hence, this indicates that the comparison between PM and HL is fair. Another relevant aspect of the coordinate distributions is the heaviness of their tails—since both PM and HL are intended to be robust methods. For this purpose, Figures 6(a)-6(d) display normal QQ plots for the coordinate distributions. These QQ plots show clear departures from a linear trend, indicating that the tails of the coordinate distributions are indeed heavy. Lastly, since the PM method is designed to take advantage of variance decay among the Fourier coefficients, the presence of this decay is confirmed Figures 7(a)-7(b).
[…]
Figure 4: Panel (a) shows log return curves for Apple (red) and Microsoft (blue), while panel (b) shows log return curves for Google (red) and Meta (blue).
[…]
Figure 5: Panels (a) and (b) respectively show histograms of second and third Fourier coefficients of $D(t)$ for (Apple, Microsoft), while panels (c) and (d) show the corresponding histograms for (Google, Meta).
[…]
Figure 6: Panels (a) and (b) respectively show normal Q-Q plots of the second and third Fourier coefficient of $D(t)$ for (Apple, Microsoft), while panels (c) and (d) show the corresponding normal Q-Q plots for (Google, Meta)
[…]
Figure 7: Panel (a) shows the estimates of the ordered values $\sigma_{(1)}\geq\cdots\geq\sigma_{(p)}$, where $\sigma_{j}^{2}$ denotes the variance of the $(j+1)$th Fourier coefficient of $D(t)$ for (Apple, Microsoft). Panel (b) shows the corresponding estimates for (Google,Meta).
@@ Appendix A Preliminaries for supplementary material
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ Y_{ij}=\varphi_{t_{j}}(X_{ij}-\mu_{j})\text{ \ \ \ and \ \ \ }M_{d}(Y)=\max_{j\in J(d)}\frac{1}{\sigma_{j}^{\tau}\sqrt{n}}\sum_{i=1}^{n}Y_{ij}-\mathbf{E}(Y_{ij}),$
@@
as recorded in [56, Eqn. 2.11]. When referring to Chebyshev’s inequality, we will typically use it in the following form for a generic random variable $U$,
@@ Appendix C Gaussian approximation
where ${\tt{r}}(R(l_{n})):=\textup{tr}(R(l_{n}))^{2}/\|R(l_{n}))\|_{F}^{2}=\frac{l_{n}^{2}}{\|R(l_{n})\|_{F}^{2}}$ is the stable rank of $R(l_{n})$. We will choose $s_{1,n}$ and $s_{2,n}$ according to
@@
$\mathbf{P}\big{(}A(s_{2,n})\big{)}\ \leq$ $\mathbf{P}(\tilde{M}_{k_{n}}(X)\leq s_{2,n})+d_{\mathrm{K}}(\mathcal{L}(M_{k_{n}}(\hat{Y})),\mathcal{L}(\tilde{M}_{k_{n}}(X))),$
@@
Consider the choice $a_{j}=\sigma_{j}^{-(1-\tau)}$ and note that under Assumption 1(iii), there is a constant $c_{0}>0$ not depending on $n$ such that the bound $a_{j}\leq c_{0}l_{n}^{(1-\tau)\beta}\!=:b$ holds for all $j\in J(l_{n})$. So, if we let $U_{j}=G_{j}(X)/\sigma_{j}^{\tau}$, then the previous two displays imply
@@
(Note that the conditions of Lemma 17 are applicable because $2\leq d_{n}\leq\frac{\omega^{2}}{4}{\tt{r}}(R(l_{n}))$ when $n$ is sufficiently large.) Furthermore, by Assumption 1(iv), we have
@@
If the conditions of Theorem 1, then there is a constant $c_{1}>0$ not depending on $n$ such that the following bound holds when $s_{1,n}=c_{1}k_{n}^{-\beta(1-\tau)/2}(2\vee\log(n))$,
@@
[…]
@@
hold simultaneously with probability at least $1-cn^{-(2+\delta)}$ Consequently, the bound
@@
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(2+\delta)}$. Combining the last several steps and Assumption 1(iii), we conclude that the bound
@@
For the first term on the right side, Nazarov’s inequality (Lemma 20) and Assumption 1(iii) imply
@@
The first term in the bound (30) is 00 with probability at least $1-\frac{ck_{n}}{n}$ by Lemma 4, and the (deterministic) third term is $\mathcal{O}(n^{-1})$ by Lemma 10.
@@
Next we will show the second factor in the bound (31) is of order $\log(n)$ with probability at least $1-\frac{c}{n}$. Let $q=2\vee\log(n)$ and observe that under Assumption 1(iii), Lemma 10 implies
@@
where the last step uses Assumption 1(i). Finally, taking a union bound over $j\in J(k_{n})$,
@@
$\mathbf{P}\Big{(}\max_{j\in J(k_{n})}\big{|}\sum_{i=1}^{n}\Delta_{ij}\big{|}\geq n^{-\frac{1}{2}}\Big{)}$ $\ \leq\ \sum_{j\in J(k_{n})}\mathbf{P}\Big{(}\big{|}\sum_{i=1}^{n}\Delta_{ij}\big{|}\geq n^{-\frac{1}{2}}\Big{)}$
@@
Nazarov’s inequality (Lemma 20) and Assumption 1(iii) imply
@@
Since $V_{1j}$ can be expressed as $\langle v,X_{1}\rangle$ for some vector $v\in\mathbb{R}^{p}$, Assumption 1(i) implies $\|V_{1j}\|_{L^{3}}^{2}\lesssim\textup{var}(V_{1j})=1$ for each $j=1,\dots,r$. Thus $\mathbf{E}\|V_{1}\|_{2}^{3}\lesssim r^{3/2}\leq k_{n}^{3/2}$, and combining with (34) completes the proof.∎
@@ Appendix D Gaussian comparison
[…]
where $I_{r}$ denotes the identity matrix of size $r\times r$. To handle the Frobenius norm, Assumption 1(i) implies
@@ Appendix E Bootstrap approximation
hold simultaneously with probability at least $1-cn^{-(2+\delta)}$ for some constant $c>0$ not depending on $n$. Based on this and $\min_{j\in J(k_{n})}\sigma_{j}^{\tau}\gtrsim k_{n}^{-\beta\tau}$ under Assumption 1(iii), the bound
@@
holds with the same probability. The terms on the right side are handled as follows. First, $\max_{j\in J(k_{n})}|\tilde{X}_{j}|$ is of order $n^{-1/2+\epsilon/2}$ with probability at least $1-cn^{-(2+\delta)}$ by Assumption 1(iii) and Lemma 12(iii). The indicators $1\{|X_{ij}|\geq\frac{t_{j}}{4}\}$ and $1\{|\tilde{X}_{j}|\geq\frac{t_{j}}{4}\}$ are 0 with probability at least $1-cn^{-2}$ due to the argument associated with the bounds in (32), as well as Lemma 12(iii) and the fact that $\sigma_{j}n^{-\frac{1}{2}+\frac{\epsilon}{2}}\lesssim\frac{t_{j}}{4}$. After taking a union bound over $i\in\{1,\dots,n\}$ and $j\in J(k_{n})$, the event
@@
Proof. We may assume without loss of generality that $k_{n}<p$, for otherwise the quantity $d_{\textup{K}}(\mathcal{L}(M_{p}(\hat{Y})),\mathcal{L}(M_{k_{n}}(\hat{Y})))$ is zero. For any $t\in\mathbb{R}$, define the events
@@
Finally, Lemma 13(iv) implies there is a constant $c>0$ not depending on $n$ that that the bound
@@
[…]
@@
Proof. We may assume without loss of generality that $\mathbf{E}(X_{1j})=0$ for all $j\in\{1,\dots,p\}$. Observe that Assumption 1(i) gives

### v2

@@ 4.2 Functional data
Preparation of stock data. Price data for 4 stocks (Apple, Nvidia, Moderna, and JP Morgan) were collected from the Alpha Vantage database (alphavantage2024) during every minute of trading between March 1, 2024 and March 22, 2024, including pre-market and after-hours trading. The data were divided into 100-minute intervals (normalized to unit length), and each interval was divided into time points $t_{0},t_{1},\ldots,t_{100}$ spaced one minute apart. For a given stock, letting $P(t_{j})$ denote its price at time $t_{j}$, we computed the cumulative log return curve within the interval as $R(t_{j})=\sum_{\ell=1}^{j}\log\big(P(t_{\ell})/P(t_{\ell-1})\big)=\log\big(P(t_{j})/P(t_{0})\big)$, $j=1,\ldots,100$. In this way, we obtained one discretely observed realization of the function $R(t)$ over each 100-minute interval. To promote independence among these functional observations, we only retained them from every other 100-minute interval, ensuring that they are separated by gaps of 100 minutes. We also excluded functional observations that were obtained from intervals with missing data.
@@
Discussion of results. In Figures 3(a)-3(d), the rejection rate curves for the proposed method are labeled by PM, and the corresponding curves based on the intervals proposed in (Fan:2023) are labeled HL. Recall that $h=0$ under the null hypothesis $\mathsf{H}_{0}$, and so the value of a curve at $h=0$ represents the empirical level. It is clear that in all four panels of Figure 3, the empirical levels of both methods closely match the nominal level of $\alpha=5\%$, marked with a dashed horizontal line.
However, the methods differ markedly in terms of power, which is represented by the values of the curves at $h>0$. In the three settings based on the stock data of Apple, Nvidia, and JP Morgan, the power of PM increases steadily with $h$, whereas the power of HL stays relatively flat. In the setting based on Moderna stock data, the two methods are more competitive, but even here, the power of PM is still noticeably higher for most values of $h$. The power advantage of PM is understandable in light of the characteristics of the simultaneous confidence intervals $\hat{\mathcal{I}}_{1},\dots,\hat{\mathcal{I}}_{p}$ for $\beta_{1},\dots,\beta_{p}$. (Recall that both methods reject the null hypothesis $\mathsf{H}_{0}\mathrel{\mathop{\ordinarycolon}}h=0$ whenever any of their associated intervals exclude 0.) As was observed in Section 4.1, the intervals produced by PM tend to be tighter than those produced by HL, and tighter intervals make it easier to exclude 0, resulting in higher power.
@@ Appendix A Preliminaries for supplementary material
[…]
@@
as recorded in [vershynin2012introduction, Eqn. 2.11]. When referring to Chebyshev’s inequality, we will typically use it in the following form for a generic random variable $U$,
@@ Appendix C Gaussian approximation
where ${\tt{r}}(R(l_{n}))\mathrel{\mathop{\ordinarycolon}}=\textup{tr}(R(l_{n}))^{2}/\|R(l_{n}))\|_{F}^{2}=\frac{l_{n}^{2}}{\|R(l_{n})\|_{F}^{2}}$ is the stable rank of $R(l_{n})$. We will choose $s_{1,n}$ and $s_{2,n}$ according to
@@
$\mathbf{P}\big(A(s_{2,n})\big)\ \leq\$$\mathbf{P}(\tilde{M}_{k_{n}}(X)\leq s_{2,n})+d_{\mathrm{K}}(\mathcal{L}(M_{k_{n}}(\hat{Y})),\mathcal{L}(\tilde{M}_{k_{n}}(X))),$
@@
Consider the choice $a_{j}=\sigma_{j}^{-(1-\tau)}$ and note that under Assumption 1((iii)), there is a constant $c_{0}>0$ not depending on $n$ such that the bound $a_{j}\leq c_{0}l_{n}^{(1-\tau)\beta}\!=\mathrel{\mathop{\ordinarycolon}}b$ holds for all $j\in J(l_{n})$. So, if we let $U_{j}=G_{j}(X)/\sigma_{j}^{\tau}$, then the previous two displays imply
@@
(Note that the conditions of Lemma 17 are applicable because $2\leq d_{n}\leq\frac{\omega^{2}}{4}{\tt{r}}(R(l_{n}))$ when $n$ is sufficiently large.) Furthermore, by Assumption 1((iv)), we have
@@
If the conditions of Theorem 1 hold, then there is a constant $c_{1}>0$ not depending on $n$ such that the following bound holds when $s_{1,n}=c_{1}k_{n}^{-\beta(1-\tau)/2}(2\vee\log(n))$,
@@
$\begin{split}\bigg\|\max_{j\in J(k_{n})^{c}}\frac{\sum_{i=1}^{n}\hat{Y}_{ij}}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}^{q}\ \leq\ \sum_{j\in J(k_{n})^{c}}\Bigg(2^{q}\bigg\|\frac{\sum_{i=1}^{n}\hat{Y}_{ij}-\mathbf{E}(\hat{Y}_{ij}|X^{\prime})}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg\|_{L^{q}|X^{\prime}}^{q}\ +\ 2^{q}\bigg|\frac{\sum_{i=1}^{n}\mathbf{E}(\hat{Y}_{ij}|X^{\prime})}{\hat{\sigma}_{j}^{\tau}\sqrt{n}}\bigg|^{q}\Bigg)\end{split}$
@@
hold simultaneously with probability at least $1-cn^{-(2+\delta)}$ . Consequently, the bound
@@
holds simultaneously over all $j\in J(k_{n})^{c}$ with probability at least $1-cn^{-(2+\delta)}$. Combining the last several steps and Assumption 1((iii)), we conclude that the bound
@@
For the first term on the right side, Nazarov’s inequality (Lemma 20) and Assumption 1((iii)) imply
@@
The first term in the bound (29) is 0 with probability at least $1-\frac{ck_{n}}{n}$ by Lemma 4, and the (deterministic) third term is $\mathcal{O}(n^{-1})$ by Lemma 10.
@@
Next we will show the second factor in the bound (30) is of order $\log(n)$ with probability at least $1-\frac{c}{n}$. Let $q=2\vee\log(n)$ and observe that under Assumption 1((iii)), Lemma 10 implies
@@
where the last step uses Assumption 1((i)). Finally, taking a union bound over $j\in J(k_{n})$,
@@
$\mathbf{P}\Big(\max_{j\in J(k_{n})}\big|\sum_{i=1}^{n}\Delta_{ij}\big|\geq n^{-\frac{1}{2}}\Big)\$$\ \leq\ \sum_{j\in J(k_{n})}\mathbf{P}\Big(\big|\sum_{i=1}^{n}\Delta_{ij}\big|\geq n^{-\frac{1}{2}}\Big)$
@@
Nazarov’s inequality (Lemma 20) and Assumption 1((iii)) imply
@@
Since $V_{1j}$ can be expressed as $\langle v,X_{1}\rangle$ for some vector $v\in\mathbb{R}^{p}$, Assumption 1((i)) implies $\|V_{1j}\|_{L^{3}}^{2}\lesssim\textup{var}(V_{1j})=1$ for each $j=1,\dots,r$. Thus $\mathbf{E}\|V_{1}\|_{2}^{3}\lesssim r^{3/2}\leq k_{n}^{3/2}$, and combining with (33) completes the proof.∎
@@ Appendix D Gaussian comparison
$\sup_{t\in\mathbb{R}}\Big|\mathbf{P}\big(M_{k_{n}}^{*}(X)\leq t\big|X\big)\ -\ \mathbf{P}\big(\tilde{M}_{k_{n}}(X)\leq t\big)\Big|\ \leq\ 2\big\|W_{r}-I_{r}\big\|_{F},$
where $I_{r}$ denotes the identity matrix of size $r\times r$. To handle the Frobenius norm, Assumption 1((i)) implies
@@ Appendix E Bootstrap approximation
hold simultaneously with probability at least $1-cn^{-(2+\delta)}$ for some constant $c>0$ not depending on $n$. Based on this and $\min_{j\in J(k_{n})}\sigma_{j}^{\tau}\gtrsim k_{n}^{-\beta\tau}$ under Assumption 1((iii)), the bound
@@
holds with the same probability. The terms on the right side are handled as follows. First, $\max_{j\in J(k_{n})}|\tilde{X}_{j}|$ is of order $n^{-1/2+\epsilon/2}$ with probability at least $1-cn^{-(2+\delta)}$ by Assumption 1((iii)) and Lemma 12(iii). The indicators $1\{|X_{ij}|\geq\frac{t_{j}}{4}\}$ and $1\{|\tilde{X}_{j}|\geq\frac{t_{j}}{4}\}$ are 0 with probability at least $1-cn^{-2}$ due to the argument associated with the bounds in (31), as well as Lemma 12(iii) and the fact that $\sigma_{j}n^{-\frac{1}{2}+\frac{\epsilon}{2}}\lesssim\frac{t_{j}}{4}$. After taking a union bound over $i\in\{1,\dots,n\}$ and $j\in J(k_{n})$, the event
@@
[…]
@@
Finally, Lemma 13(iv) implies there is a constant $c>0$ not depending on $n$ that the bound
@@
$|\mathbf{E}(\hat{Y}_{1j}|X^{\prime})|$ $\,\leq\,c\sigma_{j}^{4}\hat{\sigma}_{j}^{-3}n^{-\frac{3}{2}}\ \ \textup{ (almost surely)}$
@@
Proof. We may assume without loss of generality that $\mathbf{E}(X_{1j})=0$ for all $j\in\{1,\dots,p\}$. Observe that Assumption 1((i)) gives